### v2

# MixLM: High-Throughput and Effective LLM Ranking via Text-Embedding Mix-Interaction

Abstract.
[…]
Semantic Search, Large Language Models, Learning to Rank
copyright: acmlicensedjournalyear: 2025conference: Preprint; 2025; Preprintccs: Computing methodologies Information retrievalccs: Computing methodologies Natural language processing11footnotetext: Equal key contributors.22footnotetext: Work done while at LinkedIn.33footnotetext: Correspondence authors.
1. Introduction
Large Language Models (LLMs) have been widely adopted in industrial applications due to their strong generative and predictive capabilities. In search and recommender systems, LLMs deliver substantial relevance improvements by capturing subtle linguistic cues and conceptual relationships that traditional lexical methods often miss (Qin et al., 2024; Meng et al., 2025; Hou et al., 2024; Wu et al., 2024a). When used as relevance judges or rankers, LLMs enable semantic retrieval and ranking that better aligns with user intent.
Despite these strengths, deploying LLMs in industrial search and recommendation systems remains challenging due to their high computational cost, particularly when processing long text inputs under strict latency and throughput constraints. A common cross-encoder approach forms prompts that combine user queries with candidate item descriptions and asks the LLM to assess relevance (Qin et al., 2024; Meng et al., 2025; Ma et al., 2024; Pradeep et al., 2023). Such prompts often contain thousands of tokens, resulting in heavy prefill-dominated inference and quadratic attention costs with respect to context length. In practice, this forces systems to either replace LLMs with smaller models or drop informative features from online serving pipelines (Dey et al., 2025; Wang et al., 2024; Shang et al., 2025). Reducing input context length while preserving semantic expressiveness is therefore a central challenge for LLM-based ranking.
In this work, we introduce MixLM, an LLM-based ranking framework designed to substantially reduce input context length while retaining the semantic richness of full-text LLM rankers. MixLM compresses each candidate document—often thousands of tokens long—into a small set of learned embedding tokens using an encoder LLM. These embeddings are stored in a nearline cache and, at inference time, are mixed with the user’s natural-language query and auxiliary text and processed by a ranker LLM. By replacing long document text with a handful of embedding tokens, MixLM reduces computational complexity by orders of magnitude while preserving rich query–document interactions.
Beyond the modeling formulation, we describe the end-to-end training and deployment of MixLM in LinkedIn’s production semantic job search system. We introduce a joint training pipeline that optimizes both encoder and ranker LLMs and propose distillation losses that transfer knowledge from a strong full-text LLM ranker into the mixed-interaction architecture. We also analyze scaling behavior with respect to training data size and the number of embedding tokens per item.
From a systems perspective, we present a detailed analysis of MixLM’s impact on serving infrastructure. Context compression via embedding tokens enables substantially higher ranking throughput, which is further amplified by nearline caching and amortized prefill computation. Together, these optimizations compound to deliver large efficiency gains for mixed-input LLM serving.
We validate MixLM at industrial scale through deployment in LinkedIn’s AI-driven Job Search. Under the same latency budget, MixLM:
[…]
  • preserves the relevance quality of a strong full-text LLM baseline.
These efficiency gains enabled full-traffic deployment of LLM-based ranking and resulted in a 0.47% lift in Daily Active Users (DAU) in a large-scale online A/B experiment.
Summary of Contributions.
  • We introduce MixLM, a mixed-input LLM ranking framework that replaces long item text with compact embedding tokens to reduce prompt length.
  • We propose an end-to-end training pipeline with distillation losses that align mixed-input ranking with full-text LLM behavior.
  • We present a production-ready serving stack for mixed-input LLM ranking, achieving order-of-magnitude improvements in GPU throughput.
  • We share practical lessons from deploying MixLM in LinkedIn’s semantic job search at full production scale.
Together, these contributions demonstrate a practical and scalable path for deploying LLM-based ranking systems that combine deep semantic understanding with real-time efficiency.
2. Semantic Search
Semantic search retrieves and ranks items based on semantic similarity rather than lexical overlap. A standard pipeline consists of dense representation learning, large-scale vector retrieval (e.g., GPU-accelerated exhaustive search), high-capacity reranking to model query–item interactions, and downstream business logic such as personalization and filtering.
We focus on the ranking component and adopt a pointwise formulation in which a model consumes the query, candidate item, and optional user context to predict relevance or user engagement. Relevance labels are defined by product policy, and models are trained to optimize NDCG@10 using policy supervision.
Given a query $q$ and candidate item $i$, a cross-encoder ranker constructs a single textual prompt
(1) $\text{prompt}(q,i)=\text{system prefix},q,i,$
where the system prefix includes instructions guiding the LLM to assess semantic match.
Since relevance is treated as a binary classification task***We do not consider setwise or listwise ranking in this work., the prompt is processed by an LLM with a classification head that estimates
(2) $p_{\text{yes}}(q,i)=\mathbb{P}(i~\text{is relevant to}~q)\in[0,1],$
with higher $p_{\text{yes}}$ indicating stronger relevance.
3. Architecture of MixLM
Figure 1. Architecture of MixLM. In the offline stage, item documents are processed by an encoder LLM to obtain compact embedding representations, which are stored in a near-line cache. At serving time, these embeddings are fetched and concatenated with the user’s query and optional auxiliary text features, and the resulting mixed input is passed to the ranker LLM to generate relevance scores.
As the computational complexity of LLM inference scales with input length, reducing the number of tokens can substantially lower inference costs. In this work, we propose a text–embedding mix-interaction architecture that compresses item text description into an embedding vector using an encoder LLM. The resulting embeddings are then combined with other textual inputs and fed into a ranker LLM for output prediction. Because the item embeddings are precomputed and stored in a near-line cache, the number of tokens that must be processed during online inference is greatly reduced, leading to significant efficiency gains.
Detailed Architecture
Our ranking architecture consists of two components: an encoder LLM, which computes embeddings that summarize item information, and a ranker LLM, which predicts the final task scores. See Figure 1 for an overview.
Before diving deeper into our model architecture, let us recall that an LLM can be decomposed into three main components: The input embeddings layer, the transformer blocks and the optional output head. For an LLM, the input embeddings layer creates the input’s embedding representation. The transformer layers then form the output embedding representations, given the input embeddings. Finally, the LLM output is generated via an optional output head. For example, for a relevance classification model (in contrast to generative LLMs), we might assume the output head is a binary classification head which generates the estimated probability that the query and item are relevant.
For a sequence of length $T$ and a fixed set of tokenized vocabulary $\mathbb{V}$, we let $g_{R}(\cdot;\omega_{R}):\mathbb{V}^{T}\rightarrow\mathbb{R}^{T\times H}$ to denote the input embedding layer of the Ranker LLM, parameterized by $\omega_{R}$, with a hidden dimension of $H$. We also let $f_{R}(\cdot;\theta_{R}):\mathbb{R}^{T\times H}\rightarrow[0,1]$ denote the Ranker LLM decoder and output head, parameterized by $\theta_{R}$, where the output quantifies the probability of query-item match (i.e., $p_{\text{yes}}$). The Ranker LLM can thus be expressed as
(3) $F_{R}(\cdot;\Theta_{R})=(f_{R}\circ g_{R})(\cdot;\omega_{R},\theta_{R}):\mathbb{V}^{T}\rightarrow[0,1],$
where $\Theta_{R}=[\omega_{R},\theta_{R}]$ represents all ranker parameters.
Similarly, let $\Theta_{E}=[\omega_{E},\theta_{E}]$ denote the parameters of the Encoder LLM. The Encoder LLM for a sequence of length $T$ can be parameterized as
(4) $F_{E}(\cdot;\Theta_{E})=(f_{E}\circ g_{E})(\cdot;\omega_{E},\theta_{E}):\mathbb{V}^{T}\rightarrow\mathbb{R}^{T\times H},$
where $g_{E}(\cdot;\omega_{E}):\mathbb{V}^{T}\rightarrow\mathbb{R}^{T\times H}$ is the embedding layer and $f_{E}(\cdot;\theta_{E}):\mathbb{R}^{T\times H}\rightarrow\mathbb{R}^{T\times H}$ is the transformer that maps token embeddings to hidden representations (note that we do not have an output head for the encoder model).
In our mix-interaction system, we decompose the prompt in (1) into two parts, corresponding to ranker and encoder parts, and tokenize each part:
(5) $X_{R}$ $=\text{tokenize}([\text{system prefix},q])\in\mathbb{V}^{T_{R}}$
$X_{E}$ $=\text{tokenize}([i])\in\mathbb{V}^{T_{E}}$
where $T_{R},T_{E}$ are the lengths of $X_{R},X_{E}$. Note that $X_{E}$ does not depend on the query $q$ but only the item description $j$, and therefore, we can encode and store $X_{E}$ nearline/offline. We then let $h_{E}=f_{E}(X_{E})\in\mathbb{R}^{T_{E}\times H}$ to be the encoded item. To compress this item representation, we sample $h_{E}$ to a shorter sequence of length $T_{S}$ (specifically in MixLM, we use last N sampling), yielding $h_{S}=\mathrm{Samp}(f_{E}(X_{E}))\in\mathbb{R}^{T_{S}\times H}$. In general, one can consider several sampling techniques. In our work, for a give $T_{S}\geq 1$, we sample the $T_{S}$ embeddings in $h_{E}$ corresponding to the last $T_{S}$ input tokens. In particular, if $T_{S}=1$, we only take the embedding representation of the last input token.
Next, the input token embeddings of the prefix $X_{R}$ are computed via the ranker embedding layer, i.e., $h_{R}=g_{R}(X_{R})$. The two representations $h_{R},h_{S}$ are concatenated to form the combined embedding sequence $h_{R};h_{S};h_{EOS}]\in\mathbb{R}^{(T_{R}+T_{S}+1)\times H}$, followed by a special end of sentence token. The final “yes” probability is predicted by the ranker transformer:
(6) $p_{\text{yes}}(q,i;\Theta)=f_{R}\left([h_{R};h_{S};h_{EOS}]\right)\in[0,1],$
where $\Theta=[\Theta_{R},\Theta_{E}]$ denotes the full set of trainable parameters.
Since $h_{S}=\mathrm{Samp}(F_{E}(X_{e}))$ can be precomputed and stored in a near-line cache, the effective sequence length during online inference is reduced from to $T_{R}+T_{S}+1$. In practice, $T_{S}\ll T_{E}$, where we might take $T_{S}=1$ and $T_{E}$ is 2100 in p99, representing orders of magnitudes of reduction in the prompt length. Furthermore, when ranking multiple items, $X_{R}$ is shared across all prompts that correspond to the same user and query. As the result, the Key–Value (KV) cache of $X_{R}$ of be used leading to further reduction of computational complexity.
4. Training
In this section, we describe the models, datasets, and training pipeline used to deploy the MixLM framework in LinkedIn’s semantic job search system. MixLM requires jointly training a ranker LLM and an encoder LLM, with the key requirement that the encoder’s output embeddings align with the ranker’s input embedding space so that encoded item representations can be consumed directly by the ranker (via concatenation in Eq. (6)). As long as the ranker’s hidden size matches the encoder’s output dimension, the two models may differ architecturally. For simplicity and training efficiency, however, we use the same 0.6B-parameter architecture for both the ranker and encoder.
We adopt a three-stage training recipe for MixLM, summarized in Table 1. In Stage I, we perform supervised fine-tuning of a 0.6B pretrained model on a domain-specific dataset; the resulting checkpoint serves as the base initialization for the ranker. In Stage II, we train a pure-text ranker using full item descriptions and labeled ground truth. This model achieves high ranking quality and is treated as a teacher for subsequent distillation. In Stage III, we co-train the ranker and encoder using custom loss functions that distill knowledge from the text-based teacher into the mixed-input architecture. The ranker is initialized from the Stage I checkpoint, while the encoder is initialized from a 0.6B General Text Embedding (GTE) model trained with contrastive learning.
Table 1. MixLM Training Stages: In the stage I, we use the chain-of-thought reasoning traces from our in-house relevance judge to pretrain a 0.6B model on the domain-specific data. In stage II, we apply Supervised Fine-Tuning (SFT) on the model from stage I with the text-only ranking dataset, which serves as a teacher model in the next stage. In stage III, we co-train the encoder and ranker LLMs using a ranking dataset. We include several different loss functions in this stage, including the ranking SFT loss, a distillation loss and a self-alignment regularization.
Stages Stage I: Domain-specific Fine-Tuning Stage II: Ranking Teacher Training Stage III: Joint Encoder-Ranking Training
Data Reasoning Dataset Ranking Dataset Ranking Dataset
Number of Samples 180k 10.9M 10.9M
Input Prompt Type Text (chain-of-thoughts) Text Text + Embedding
Maximum Sequence Length 2 048 2 048 2 176
Trainable Module(s) Pretrained LLM Ranker LLM Encoder LLM & Ranker LLM
Training Objective Distillation from relevance judge SFT SFT, Distillation from ranking teacher, self-alignment
Figure 2. Detailed training pipeline of training state III.
4.1. Datasets
4.1.1. Label Generation
We sampled query–item pairs from real user logs, and train a 7B LLM as our relevance judge to generate 5-point graded scores along several dimensions, together with rationales, such that it aligns with the product policy. The integer labels were subsequently normalized to continuous values in the range $[0,1]$ to create the ground truth labels $p^{*}_{\text{yes}}(q,j)$ for query-item pairs.
This 7B in-house LLM operates using the following template:
[CRITERIA]   : <matching guidelines>
[EXAMPLES]   : <reasoning and output format>
[QUERY]      : <query text>
[CANDIDATE]  : <item text>
Analyze the query--candidate pair and determine whether
they are a good match. Return a single matching score
(0/1/2/3/4) with detailed reasoning.
We gather model’s chain-of-thought traces (Wei et al., 2022), as well as the final matching score which we use in different stages of training.
4.1.2. Teacher Training Prompts in Stage II
For the textual teacher training, we applied the following chat template to the ranker LLM:
[SYSTEM]     : Output only ’Yes’ or ’No’ based on how well
               the item matches the query; no extra text.
[QUERY]      : <query text>
[CANDIDATE]  : <item text>
yes or no ?
4.1.3. Mixed Input Ranking Training Prompts in Stage III
For the joint ranking training (co-train the encoder LLM and ranker LLM), we applied the following chat template to the ranker LLM:
[SYSTEM]     : Output only ’Yes’ or ’No’ based on how well
               the item matches the query; no extra text.
[QUERY]      : <query text>
[CANDIDATE]  : <item embeddings> <EOS>
where the <item embeddings> is the output from encoder LLM. The encoder LLM use the following prompt “Item information: <item text>”. Note that the ranking prompt used here is significantly shorter compared to the teacher one in Section 4.1.2 which uses full item text.
4.2. Three-stage Training
Next, we provide the details of our training pipeline.
4.2.1. Stage I: Domain Reasoning Fine-Tuning
In the first stage, we train a general pretrained LLM using a domain-specific reasoning dataset to strengthen its capability for logical inference and contextual understanding over query–item pairs. This intermediate training phase enables the model to grasp the rationale behind assigning relevance scores, thereby enhancing its generalization performance in downstream ranking tasks.
Concretely, we distill (Hinton, 2014) our 7B relevance judge into a 0.6B pretrained model, using KL divergence between the logits of the two models. As for the data, we randomly select 180180K samples following the prompt given in Section 4.1.1, together with the chain-of-thoughts responses of the relevance judge model.
4.2.2. Stage II: Ranking Teacher Training
In the next step, we train a purely textual ranking model. Although this intermediate model cannot be served online due to the long input prompts, it will serve as a teacher model when training the mixed input MixLM model. As we demonstrate, using such a ranking teacher model results in better convergence and training quality.
Specifically, starting from the checkpoint available from our first training stage, we perform Supervised Fine-Tuning (SFT) using the input prompt discussed in Section 4.1.2. We generate 10.910.9M examples based on the prompt provided above, and normalize the integer matching scores, available from the 7B judge, into ground truth probabilities $(p^{*}_{\text{yes}},p^{*}_{\text{no}})$. The SFT is then performed using a KL divergence loss. We use $(\hat{p}_{\text{yes}},\hat{p}_{\text{no}})$ to denote the output probability distribution from the final checkpoint for this training stage.
4.2.3. Stage III: Joint Encoder-Ranking Training
In the final training stage, we train both ranker and encoder models to estimate relevance matching scores, using the mixed input prompt from Section 4.1.3. Specifically, the relevance scores are calculated using (6). The trainable parameters are $\Theta=[\Theta_{R},\Theta_{E}]$ which encompasses all parameters of ranker and encoder models.
To this end, we initialize the ranker module in Figure 1 via the final checkpoint from our first stage training. For the encoder module, we use a pretrained 0.6B GTE model. We use the same 10.9M sample dataset from the previous stage for training. Given the specific structure of the ranking stack, we have observed using custom loss functions can improve the training quality and the final models’ ranking quality. Below we discuss the details of the loss functions we use.
Soft-Label SFT Loss
Our first loss function is a KL divergence SFT loss, that ensures the predicted probabilities from MixLM are close to the ground truth. This loss function is given as
$\mathcal{L}_{\text{SFT}}(\Theta)=\text{KL}\left((p^{*}_{\text{yes}},p^{*}_{\text{no}})||({p}_{\text{yes}}(\Theta),{p}_{\text{no}}(\Theta))\right)$
where $({p}_{\text{yes}},{p}_{\text{no}})$ are the predicted probability of relevance match (see (6)), and $(p^{*}_{\text{yes}},p^{*}_{\text{no}})$ are ground truth probabilities.
Ranking Distillation Loss
In addition to comparing to the ground truth, we compare the output distribution from MixLM to the one from the ranking teacher model from Stage II. More specifically, we define
$\mathcal{L}_{\text{distill}}(\Theta)=\text{KL}\left((\hat{p}_{\text{yes}},\hat{p}_{\text{no}})||({p}_{\text{yes}}(\Theta),{p}_{\text{no}}(\Theta))\right)$
where we recall that, $(\hat{p}_{\text{yes}},\hat{p}_{\text{no}})$ is the output distribution of ranking teacher with full text prompt in Section 4.1.2.
This loss function is inspired by knowledge distillation (Hinton, 2014), where knowledge from a (typically larger) teacher model is transferred to a (usually smaller) student model. In contrast, our student model, MixLM, which consists of an encoder and a ranker, can in fact be larger than the teacher model. As discussed in Section 6, the teacher operating on pure text is easier to train to provide high-quality predictions that are cleaner than the original dataset labels. When training MixLM, the teacher label helps reducing gradient variance and facilitates better convergence.
Self-Alignment Loss
Finally, we use two self-alignment loss functions, that help to ensure the encoder model’s output embeddings are aligned with the input embeddings of the ranker model. From our experience, including these loss functions has a regularizaiton effect that can prevent overfitting. To this end, we pass the full text prompt (from Section 4.1.2) through the mixed input model’s ranker module (the model parameterized by $\Theta_{R}$ from section 4), and collect the output probability distribution (which we denote by $(\bar{p}_{\text{yes}},\bar{p}_{\text{no}})$). Additionally, we collect the last layer’s (before the output head) hidden states corresponding to the last input token, which we denote by $\bar{h}_{\text{last}}\in\mathbb{R}^{H}$. Note that these quantities are functions of only $\Theta_{R}$. For the same sample, we also collect the output probability distribution (denoted as $(p_{\text{yes}},p_{\text{no}})$) and the last layer’s hidden states for the last input token (denoted as $h_{\text{last}}\in\mathbb{R}^{H}$) when we use the mixed input prompt from Section 4.1.3. These quantities, as expected, are functions of all trainable parameters $\Theta$. Then, we use loss functions that encourage the resulting quantities from these two setups to be similar. First, we use a loss function $\mathcal{L}_{\text{hidden-align}}$ to encourage the last hidden states from the mixed input and textual prompts to be similar:
$\mathcal{L}_{\text{hidden-align}}(\Theta)=1-\mathrm{cossim}\big(h_{\text{last}}(\Theta),\bar{h}_{\text{last}}(\Theta_{R})\big)$
where $\mathrm{cossim}$ denotes the cosine similarity of two vectors. In addition, we use another alignment loss to ensure the predicted probabilities remain similar:
$\mathcal{L}_{\text{pred-align}}(\Theta)=\text{KL}\left((\bar{p}_{\text{yes}}(\Theta_{R}),\bar{p}_{\text{no}}(\Theta_{R}))||(p_{\text{yes}}(\Theta),p_{\text{no}}(\Theta))\right).$
The overall alignment loss is a linear combination of hidden state and prediction alignment losses:
$\mathcal{L}_{\text{align}}(\Theta)=\alpha\cdot\mathcal{L}_{\text{pred-align}}(\Theta)\;+\;\beta\cdot\mathcal{L}_{\text{hidden-align}}(\Theta)$
where $\alpha,\beta\geq 0$ are given hyper-parameters.
Finally, the full training objective integrates all components discussed above. Specifically, we use
$\mathcal{L}_{\text{total}}(\Theta)=\mathcal{L}_{\text{SFT}}(\Theta)+\lambda_{\text{distill}}\mathcal{L}_{\text{distill}}(\Theta)+\lambda_{\text{align}}\mathcal{L}_{\text{align}}(\Theta),$
where $\lambda_{\text{distill}}\geq 0$ and $\lambda_{\text{align}}\geq 0$ are hyper-parameters that are given a priori. Figure 2 summarizes the training pipeline for this stage.
To train MixLM more effectively, we apply a phased curriculum learning strategy, decomposing the joint optimization into two distinct phases with different loss weights, i.e., ($\lambda_{\text{distill}}$ and $\lambda_{\text{align}}$): In Phase 1, we prioritizes the space-alignment by increasing $\lambda_{\text{align}}$ and decreasing $\lambda_{\text{distill}}$, so that we can train the encoder to produce embeddings that integrate seamlessly with the ranker’s internal representations. In Phase 2, decrease $\lambda_{\text{align}}$ and increase $\lambda_{\text{distill}}$, focusing on adjusting ranker LLM to improve the relevance score prediction.
5. Inference Engine Optimization
5.1. Text–Embedding Mixed Input
To support MixLM’s mixed-interaction design, we extend the serving interface to accept precomputed embedding tensors alongside text. Embeddings are transmitted via a structured feature payload added to the gRPC request schema, consisting of a binary tensor buffer and metadata specifying shape, data type, and optional serialization format (e.g., NumPy (Harris et al., 2020)).
On the server side, the preprocessing module decodes and validates feature payloads into native tensors using metadata-driven reshaping and casting, and attaches them to the model input context. This enables direct reuse of embeddings produced by upstream pipelines, avoiding redundant computation and reducing latency. The schema remains fully backward-compatible with text-only requests and introduces minimal serialization overhead.
Table 2. Attention and linear-layer compute under different conditions, in terms of $(T_{q},T_{i},N_{i})$, where $T_{i}$ denotes full item-text length and $K$ is the MixLM compression factor (item length becomes $T_{i}/K$). $\propto$ denotes proportionality up to architecture-dependent constants. Bottom blocks instantiate $N_{i}=250$, $K=450$.
Conditions Total attention cost Total linear-layer cost
Naive (full-text items) $\propto N_{i}(T_{q}+T_{i})^{2}$ $\propto N_{i}(T_{q}+T_{i})$
Amortized prefill (full-text items) $\propto T_{q}^{2}+N_{i}(2T_{i}T_{q}+T_{i}^{2})$ $\propto T_{q}+N_{i}T_{i}$
Amortized prefill ++ MixLM $\propto T_{q}^{2}+N_{i}\bigl(2(T_{i}/K)T_{q}+(T_{i}/K)^{2}\bigr)$ $\propto T_{q}+N_{i}T_{i}/K$
Naive, $N_{i}=250$ $\propto 250(T_{q}+T_{i})^{2}$ $\propto 250(T_{q}+T_{i})$
Amortized, $N_{i}=250$ $T_{q}^{2}+500T_{i}T_{q}+250T_{i}^{2}$ $\propto T_{q}+250T_{i}$
Amortized ++ MixLM $N_{i}=250$, $K=450$ $T_{q}^{2}+\frac{10}{9}T_{i}T_{q}+\frac{1}{810}T_{i}^{2}$ $T_{q}+\frac{5}{9}T_{i}$
5.2. Shared-Prefix Prefill Optimization
In large-scale ranking, each request typically scores hundreds to thousands of items sharing the same query and user context. After MixLM compresses each item into a small number of embedding tokens (from a median of $\sim$900 tokens to two tokens in production), inference cost becomes dominated by repeated query-prefix computation.
To improve GPU utilization, we batch candidate-item prompts per query and reuse the KV cache across prompts with a shared prefix. This is particularly effective under MixLM, where item tokens are minimal and query-prefix computation dominates cost.
Let $T_{q}$ be the query-prefix length, $T_{i}$ the full-text item length, and $N_{i}$ the ranking depth. Ignoring architecture-dependent constants, attention and linear costs scale as
$\text{Attention}\propto L^{2},\qquad\text{Linear}\propto L.$
Naïve prefill (full-text).
Each item is processed independently with length $(T_{q}+T_{i})$:
$F_{\text{att}}^{\text{naive}}\propto N_{i}(T_{q}+T_{i})^{2},\qquad F_{\text{lin}}^{\text{naive}}\propto N_{i}(T_{q}+T_{i}).$
Amortized prefill (full-text).
Query-prefix attention is computed once and reused:
$F_{\text{att}}^{\text{amort, full}}\propto T_{q}^{2}+N_{i}(2T_{i}T_{q}+T_{i}^{2}),\qquad F_{\text{lin}}^{\text{amort, full}}\propto T_{q}+N_{i}T_{i}.$
Amortized prefill with MixLM.
With item compression factor $K$, each item contributes $T_{i}/K$ tokens:
$F_{\text{att}}^{\text{amort, MixLM}}\propto T_{q}^{2}+N_{i}\bigl(2(T_{i}/K)T_{q}+(T_{i}/K)^{2}\bigr),$
$F_{\text{lin}}^{\text{amort, MixLM}}\propto T_{q}+N_{i}T_{i}/K.$
Thus, item-side attention and linear costs are reduced by approximately $K$.
We implement amortization using two mechanisms:
  • In-batch prefix caching: We reuse the key–value (KV) states from the first prompt so that all items in the batch share the same query-prefix computation in a single forward pass. The attention computation merges two operations: (1)suffix tokens attend to all prefix tokens from the first prompt via dense paged attention; (2)suffix tokens attend to themselves via regular causal attention.
  • Multi-item scoring: Multiple items are concatenated into a single sequence separated by delimiters (Sundara Raman Ramachandran, 2025). FlashInfer applies item-aware masking to prevent cross-item attention.
Because MixLM reduces $T_{i}$ to only a few tokens, shared-prefix amortization eliminates most redundant computation and delivers substantial inference-throughput gains in production.
5.3. Inference Engine CPU Overhead Reduction
5.3.1. Multi-Process gRPC
After optimizing the model and the SGLang inference engine, request preprocessing in the Python gRPC service emerged as the dominant bottleneck. Although the gRPC server uses asynchronous Python, request deserialization, feature decoding, and context construction are constrained by the Global Interpreter Lock (GIL) and the single-threaded asyncio event loop. Under high request rates, event-loop contention and per-request overhead limit throughput despite asynchronous I/O.
To address this, we adopt a multiprocessing server architecture that decouples the gRPC frontend from the SGLang engine. Request handling and preprocessing are parallelized across multiple CPU processes, while the SGLang engine runs independently. This design fully utilizes multi-core CPUs and improves throughput by 40%.
5.3.2. Batch Send
To fully exploit amortized prefill, the SGLang scheduler must admit prompts sharing the same query as a single batch. We therefore introduce an explicit batch send mechanism, where all item prompts for a query are serialized and transmitted as a single ZMQ message. This ensures that the scheduler processes the batch atomically and can reliably apply in-batch prefix caching.
5.3.3. Multi-Process Parallelization
While GPU prefill executes efficiently, request scheduling on a single CPU core became a secondary bottleneck. We mitigate this by running multiple scheduler processes (e.g., five) per GPU, each assigned a dedicated GPU memory partition (e.g., 20% per worker). This parallelizes CPU-side scheduling while maintaining high GPU utilization, effectively pipelining batch preparation with GPU execution.
5.3.4. Eliminating Redundant KV-Cache Operations
SGLang’s default KV-cache behavior is designed for autoregressive decoding, which is not used in our prefill-only scoring workload. We modify cache management to immediately release KV cache memory after request completion, eliminating unnecessary allocation, eviction, and pool management. This both reduces CPU overhead on the critical path and increases effective GPU memory capacity, enabling larger batch sizes and higher concurrent throughput.
6. Experimental Results
We conduct both offline and online experiments to evaluate the ranking quality and serving efficiency of MixLM. Offline evaluation uses a held-out test set labeled by our internal LLM relevance judge. We compare against several baselines: (1) an embedding-based bi-encoder retrieval model distilled from the LLM judge (the first-stage retriever in our stack), (2) a pure-text 0.6B LLM ranker from the second training stage, representing an upper-bound model that is impractical to deploy, and (3) LinkedIn’s limited-scale production baseline that applies near-line text summarization and aggressive ranker pruning (Behdin et al., 2025).
We report NDCG@10 for ranking quality and maximum throughput measured as items scored per GPU per second, under a fixed p99 end-to-end latency constraint of 500 ms.
6.1. Efficiency and Relevance
Table 3. Ranking quality comparison of several ranking approaches. We report NDCG@10 with respect to labels provided by our relevance judge.
Model NDCG@10
Full Text 0.9432
Summarized, Pruned 0.9218
MixLM 0.9239
Embedding Retrieval 0.8380
Table 4. GPU efficiency comparison of ranking approaches. Throughput is measured as maximum items scored per GPU per second under a 500 ms latency budget.
Model QPS (Items/s/GPU) Latency (ms) GPU
Full Text 290 <<500 H100
Summarized, Pruned 2,200 <<500 H100
MixLM 22,000 <<500 H100
Embedding Retrieval $1.6\times 10^{9}$ <<100 A100
Table 3 shows that embedding-only retrieval delivers the lowest ranking quality and is therefore suitable only as a coarse first-stage filter. Full-text LLM ranking achieves the highest NDCG@10 (0.9432) but is prohibitively expensive, scoring only 290 items/s/GPU. Summarized-text LLM ranking with pruning improves throughput to 2,200 items/s/GPU while sacrificing some quality (NDCG@10 = 0.9218).
[…]
6.2. Online A/B Test
Table 5. Online A/B Test Results
[…]
Classic Job Search
Semantic Job Search (MixLM) +0.47%
Table 5 summarizes the findings. After deploying MixLM in LinkedIn’s AI Job Search and running extensive A/B tests, we observed: MixLM has on-par relevance metrics comparing with the existing limited-scale production baseline(summarized text LLM)  (Behdin et al., 2025). With the 10.0x gain in serving throughput under the same latency budget, MixLM empowered the full-traffic deployment of LLM-ranking-based Semantic Job Search and for the first time, consequently producing a significant 0.47% increase in Daily Active Users, compared with the traditional job search.
6.3. Ablation Analysis
To understand the contribution of individual components, we perform ablation studies on a smaller dataset that mirrors the distribution of the full training corpus.
6.3.1. Scaling Up
Table 6. Ablation on Dataset Size
[…]
160K
400K +0.0250
800K +0.0280
1.08M +0.0334
Table 7. Embedding Tokens per Item
[…]
1
5 +0.0017
10 +0.0044
20 +0.0111
30 +0.0158
40 +0.0172
50 +0.0198
Increasing both training data size and embedding granularity consistently improves ranking quality. For production, we retain a single embedding token per item to meet latency constraints.
6.3.2. Domain Reasoning Fine-Tuning
Table 8. Effect of Domain Reasoning Fine-Tuning
[…]
Vanilla LLM
Domain-Reasoning Tuned LLM +0.0185
Domain reasoning fine-tuning substantially improves ranking quality and is adopted as the first stage of our training pipeline.
6.3.3. Self-Alignment and Auxiliary Losses
Table 9. Ablation on Auxiliary Losses
[…]
No auxiliary loss
+self-alignment +0.0014
+distillation +0.0091
+self-alignment + distillation +0.0108
Distillation provides the dominant gain, while self-alignment further improves results when combined with distillation.
6.3.4. Inference Optimizations
Table 10. Ablation on Inference Optimizations (<<500 ms latency)
Configuration Prefix Optimization QPS (Items/s/GPU)
Raw-text None 270
Mult-Item Scoring 275
In-Batch Prefix Cache 290
Summarized, Pruned None 1,650
Multi-Item Scoring 2,100
In-Batch Prefix Cache 2,200
MixLM None 3,000
Multi-Item Scoring 20,000
In-Batch Prefix Cache 22,000
MixLM combined with in-batch prefix caching yields the highest throughput, confirming the importance of both input compression and shared-prefix amortization for high-QPS LLM ranking.
7. Related Works
7.1. LLMs for Semantic Search
Extensive prior work studies the use of LLMs for semantic search. A widely adopted and computationally efficient approach is the bi-encoder paradigm (Ma et al., 2024; Karpukhin et al., 2020), which independently embeds queries and documents into a shared vector space and is commonly deployed in production (Huang et al., 2020). However, limited query–document interaction can miss fine-grained semantics, motivating late-interaction models that introduce token-level matching (Khattab and Zaharia, 2020).
Cross-encoder models jointly encode queries and documents (or document lists), enabling richer interaction modeling. LLM-based cross-encoders achieve strong ranking performance (Sachan et al., 2022; Zhuang et al., 2024; Zhang et al., 2024; Pradeep et al., 2023; Meng et al., 2025), but their computational cost makes large-scale online serving challenging (Wang et al., 2024; Gao et al., 2020; Hofstätter et al., 2020; Xia et al., 2025).
To address these limitations, prior work leverages LLMs indirectly. The LLM-as-judge paradigm uses LLMs to generate graded relevance labels and rationales to supervise efficient student models (Wang et al., 2024; Ye et al., 2025; Tang et al., 2019; Reddi et al., 2021; Qin et al., 2023; Xia et al., 2025), while LLM-as-augmenter approaches enrich training data via paraphrasing, summarization, or hard-negative generation (Behdin et al., 2025; Li et al., 2024). Our work builds on these directions while enabling direct deployment of LLM-based ranking through model–infrastructure co-design.
7.2. Mixed-Input LLMs for Recommendation Systems
Mixed-input LLMs that combine text tokens with embedding inputs have been explored in recommendation settings. Prior work replaces parts of text prompts with embeddings from traditional models (Zhang et al., 2025b) or represents user interaction histories using embeddings rather than text (Chen et al., 2024). These methods primarily target recommendation or offline feature generation.
In contrast, we focus on relevance ranking, where nuanced semantic understanding is critical. We present a complete training recipe for mixed-input ranking, including custom losses and distillation from pure-text LLMs, and deploy the model directly for online inference with reported A/B test results. We further describe serving-system optimizations required to scale mixed-input ranking models in production.
Concurrent work by Lin et al. (2025) studies mixed-input LLMs for retrieval-augmented generation, a setting distinct from ranking and evaluated primarily offline.
7.3. Multimodal Large Language Model Training
Multimodal LLMs commonly extend pretrained text models by injecting modality-specific embeddings via lightweight adapters. Early systems such as Flamingo (Alayrac et al., 2022) and Kosmos-1 (Huang et al., 2023) treat non-text inputs as additional tokens to enable strong vision–language performance.
Later work emphasizes minimal architectural changes by aligning pretrained encoders with LLM embedding spaces. BLIP-2 (Li et al., 2023) uses a querying transformer to align vision features, while LLaVA-style models (Liu et al., 2023) project CLIP features directly into the LLM token space. MixLM follows this paradigm by injecting learned embedding tokens to replace long text inputs, adapting multimodal techniques to high-throughput semantic ranking.
8. Conclusion
In this work, we introduced MixLM, a novel framework that bridges the gap between the semantic richness of large language model–based rankers and the stringent efficiency requirements of industrial search and recommender systems. By representing candidate items as compact embedding tokens and introducing the mix-interaction mechanism, MixLM effectively reduces the input context length while maintaining the expressiveness of full-text LLMs. Our co-training strategy for online rankers and near-line embedding flows further enables tight coupling between semantic understanding and serving efficiency.
Empirical results from deployment in a large-scale professional social network demonstrate that MixLM delivers substantial system gains—10.0x times thorughput compared with summarized-text LLM, 75.9x times thorughput compared with full-text LLM, under same latency budget. MixLM enabled LLM search system’s deployment on all LinkedIn users’ traffic, thereby producing a measurable 0.47% increase in daily active users (DAU). These findings highlight the potential of mixed-input architectures to reshape the design space of LLM-based ranking systems, offering a practical path toward scalable, low-latency, and semantically powerful search in real-world environments.
Future work will explore adapting co-trained embeddings for modeling member history and exploring MixLM for retrieval tasks.